OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3. −91 −33 −25 35

OpenStudy (anonymous):

f(g(x))=2(-5x+2)+1; f(g(x))=-10x+5; f(3)=-10(3)+5=-25